### v1

@@ Abstract
Source-Free Unsupervised Domain Adaptation (SF-UDA) aims to transfer a model’s performance from a labeled source domain to an unlabeled target domain without direct access to source samples, addressing data privacy issues. However, most existing SF-UDA approaches assume the availability of abundant source domain samples, which is often impractical due to the high cost of data annotation. In this paper, we explore a more challenging scenario where direct access to source domain samples is restricted, and the source domain contains only a few samples. To tackle the dual challenges of limited source data and privacy concerns, we introduce a data-efficient, CLIP-powered dual-branch network (CDBN in short). We design a cross-modal dual-branch network that integrates source domain class semantics into the unsupervised fine-tuning of the target domain. It preserves the class information from the source domain while enhancing the model’s generalization to the target domain. Additionally, we propose an unsupervised optimization strategy driven by accurate classification and diversity, which aims to retain the classification capability learned from the source domain while producing more confident and diverse predictions in the target domain. Extensive experiments across 31 transfer tasks on 7 public datasets demonstrate that our approach achieves state-of-the-art performance compared to existing methods.
@@
Source-Free, Unsupervised Domain Adaptation, Prompt Learning, CLIP-powered, Data-Efficient
@@ I Introduction
Deep learning has achieved remarkable performance in various computer vision tasks [3, 58, 48]. However, these achievements largely rely on the assumption that the training and test data are independently and identically distributed (i.i.d.) [19]. In real-world scenarios, differences in data collection conditions, environmental factors, and other variables often cause domain shifts between the training data (source domain) and test data (target domain), leading to a significant reduction in model performance. To address this issue, previous works [24, 7] have proposed unsupervised domain adaptation (UDA), employing adversarial training [43, 7], metric learning [24], and other techniques to learn domain-invariant features, thereby facilitating the transfer of models trained on the source domain to the unlabeled target domain.
@@
Despite the effectiveness of UDA methods in addressing domain shifts, they typically assume that the source domain is accessible during target domain training and that it contains sufficient labeled data for training. However, this assumption is often unrealistic in practical settings due to the high cost of data annotation and the need to protect data privacy. For instance, in the diabetic retinopathy dataset used in clinical research [11], each image is annotated by 7 to 8 board-certified ophthalmologists, significantly increasing the cost of labeling [53]. Additionally, in fields such as medicine and remote sensing [35], labeled data often cannot be distributed or reused due to privacy protection and commercial copyright concerns.
@@
Figure 1: Scheme 1 illustrates the issue with previous SF-UDA methods, where using a fully trained model from a source domain as an information source for a target domain leads to poor performance when source domain samples are limited. To address this issue, Scheme 2 presents our proposed solution, which involves transferring category feature representations obtained from a source-domain-based CLIP prompt learning method to the target domain, effectively resolving the SF-UDA problem in scenarios with few source domain samples.
@@
In response to the dual challenges of inaccessible source domain data and limited labeled samples, researchers have developed two types of methods. Source-free unsupervised domain adaptation (SF-UDA) methods [23, 43, 44] address the issue of inaccessible source data by enabling model transfer to the target domain without requiring direct access to the source domain. Meanwhile, few-shot unsupervised domain adaptation (FS-UDA) methods [53, 17, 22] tackle the problem of limited labeled data by facilitating effective model transfer with minimal source samples. However, methods that address the dual challenges of limited access to source domain data and small amounts of source domain samples, known as SF-UD methods under few-shot source domain scenarios (named in FS-SF-UDA in this paper), have yet to be fully explored.
@@
In recent years, prompt learning for CLIP models [16, 56, 9] has demonstrated significant advantages in few-shot classification tasks. These methods typically leverage CLIP as a backbone network and employ prompt learning techniques to achieve strong downstream task performance, even with only a small number of samples per class. Inspired by this, we propose a data-efficient CLIP-powered Dual-Branch Network, (dubbed as CDBN) for FS-SF-UDA.
@@
With only a few labeled source domain samples, our goal is to efficiently extract category-specific semantic information from the source domain and transfer it to the target domain as a reference to assist in the unsupervised training process. Specifically, as shown in Scheme 1 of Figure 1, traditional supervised training methods perform poorly when source domain samples are limited. Therefore, as illustrated in Scheme 2 of Figure 1, we extract generalized category text features from the small labeled source domain samples using prompt learning. These text features serve as carriers of source domain category semantics, which are then transferred to the target domain to guide unsupervised training.
@@
However, the limited number of source domain samples, combined with the domain gap, poses a challenge as it introduces noise and domain shift into the transferred category text features. To address this, we designed a cross-modal dual-branch network consisting of a cross-domain feature transfer branch and a target-specific feature learning branch. The cross-domain feature transfer branch utilizes high-confidence samples from the target domain to transfer the learned source domain category text features, reducing the impact of domain discrepancy. Meanwhile, the target-specific feature learning branch focuses on learning independent category features from the target domain, avoiding over-reliance on noisy source domain text features. The synergy between these two branches effectively enhances the feature representation capabilities in the target domain.
@@
Furthermore, since there are no labels available during target domain training, we propose an unsupervised optimization strategy driven by accurate classification and diversity. This strategy includes three core loss functions: first, a cross-entropy loss based on high-confidence samples to ensure accurate classification in the target domain; second, a multi-view consistency loss to encourage the model to extract consistent feature representations from different perspectives, enhancing generalization; and third, a mutual information maximization loss between inputs and outputs to encourage diversity in the category distribution, preventing overfitting to simple classes. Together, these three components ensure that the model can accurately and generatively learn target domain features under unsupervised conditions.
@@
We conducted experiments on 31 transfer tasks across 7 public domain adaptation datasets. Despite the dual challenges of limited source samples and restricted access to source domain data, CDBN achieved state-of-the-art performance on multiple datasets and showed competitive results comparable to the best existing methods on others. These outcomes demonstrate the robustness and effectiveness of our CDBN approach in addressing the FS-SF-UDA problem.
@@
  • We introduce a cross-modal dual-branch network structure that effectively leverages source domain category semantic information during the unsupervised fine-tuning process in the target domain, while mitigating the effects of source domain noise and domain discrepancy. Additionally, it combines the strengths of both branches, significantly enhancing performance in the target domain.
@@ II-B Few-Shot Unsupervised Domain Adaptation
Few-Shot Unsupervised Domain Adaptation (FS-UDA) aims to address the challenge of having limited labeled samples in the source domain during training on the target domain, even though source domain samples are accessible. Previous methods have explored different scenarios to tackle this issue. For instance, Yue et al. [53] proposes a method to address the problem of label sparsity in the source domain. In situations where a large number of source domain samples are available but only a small portion are labeled, they leveraged cross-domain self-supervised learning to adapt the model to the target domain. Similarly, Wang et al. [47] presents a method to handle the long-tailed distribution of source domain data. In cases where only a few samples are available for certain hard-to-access categories in the source domain, they improved model accuracy through generative feature augmentation. Furthermore, Li et al. [22] proposes a Multi-Prototype Alignment framework to transfer the model to the target domain when only a few samples per category are available in the source domain. This approach effectively addresses the scenario where there are very few samples for each category. In this paper, we tackle the challenges present in both FS-UDA and SF-UDA by training a high-performing model on the target domain, despite having only a few samples per class in the source domain and no direct access to source domain samples during target domain training.
@@ II-C Prompt Learning for CLIP
In recent years, vision-language pretraining models have made significant progress in image representation learning by leveraging supervision from natural language to interpret images. Among these models, the Contrastive Language-Image Pre-training (CLIP [30]) model is particularly representative. CLIP consists of a text encoder and an image encoder, pretrained on 400 million image-text pairs collected from the internet, successfully aligning the feature spaces of both modalities. This approach demonstrates strong generalization capabilities in downstream tasks and enables zero-shot transfer in image classification tasks through manually specified text prompts. Inspired by prompt learning techniques in natural language processing, Zhou et al. [56] introduces a learnable context optimization technique to adapt vision-language models like CLIP to image recognition tasks. Instead of relying on manually designed prompts, CoOp (Context Optimization) uses learnable vectors to model contextual words while keeping the pretrained parameters frozen. This method allows CoOp to exhibit superior performance with limited labeled data, showing a significant improvement over manual prompt engineering in few-shot scenarios. To further enhance the alignment between visual and textual representations, Khattak et al. [16] proposes a method that applies prompt learning to both the image and text branches simultaneously, fostering a strong coupling between visual and language prompts to ensure their mutual synergy. As a result, Maple demonstrates significantly better generalization to novel classes, surpassing CoOp in performance.
Figure 2: The framework of CDBN, where $E_{img}$ represents the image encoder, $E_{text}$ represents the text encoder, $T_{t}$ denotes the class-shared soft prompt, and $V_{c}^{L}$ indicates the learnable token for the substitute class name of class c. Solid lines represent data flow, while dashed lines indicate the flow of values without gradient backpropagation. The flame symbol represents gradient updates, and the snowflake symbol indicates frozen gradients.
@@ III-A Problem Definition
Source-free unsupervised domain adaptation (SF-UDA) is a important sub-problem of unsupervised domain adaptation, focused on scenarios where source domain samples cannot be directly accessed during the target domain training. This paper further explores an even more challenging scenario: the source domain samples are not only inaccessible but also limited in number.
Specifically, consider a labeled source domain dataset with only a small number of samples,$D_{S}=\{(x_{s}^{i},y_{s}^{i})\}_{i=1}^{n_{s}}$,and an unlabeled target domain dataset $D_{T}={x_{t}^{i}}{i=1}^{n_{t}}$. $D_{S}$ and $D_{T}$ share the same labels but have different feature distributions and cannot be trained simultaneously. The process involves first training a model $M_{S}$ with $D_{S}$, then using $M_{S}$ and $D_{T}$ for further training, and finally evaluating the model’s performance on $D_{T}$.
@@ III-B Overview of CDBN
In this work, we aim to transfer the performance of a model trained with limited supervised samples from the source domain to unlabeled samples in the target domain. As illustrated in Figure 2, we propose a CLIP-Powered dual-branch network comprising a cross-domain feature transfer branch and a target-specific feature learning branch.
The cross-domain feature transfer branch seeks to enhance the class textual features $G^{S}$ learned from the source domain by leveraging high-confidence target domain image features $W_{e}$ through a cross-attention layer, thereby transferring these class textual features to the target domain. Specifically, after supervised training on the source domain, we obtain the class textual features by generating textual feature representations for each class using CLIP’s text encoder $E_{img}$. High-confidence target domain samples $D_{T}^{\prime}$ are selected by computing the cosine similarity between the source domain’s learned class feature representations and target domain samples, selecting the top K most relevant samples for each class. The target-specific feature learning branch aims to capture the class semantic information of the target domain by learning target-specific class textual features through unsupervised learning, utilizing a class-shared soft prompt. Subsequently, we feed the class textual features obtained from both branches into a fusion logits network(FLN), which merges the logits from both branches to produce the model’s final prediction results.
Additionally, Given the lack of labels in the target domain, we propose an unsupervised optimization strategy that balances class accuracy and prediction diversity. First, we employ a cross-entropy loss $L_{CE}^{T}$ on high-confidence samples filtered by the source domain’s class textual features to retain their original classification capability. Second, we introduce a multi-view consistency loss $L_{cons}$ to encourage more target domain samples to participate in model training, further enhancing the model’s accuracy. Finally, by computing the mutual information loss $L_{IM}$ between the model’s output class distribution and the input data, we encourage the model to achieve deterministic results while maintaining prediction diversity, preventing it from collapsing into trivial class biases.
III-C Integrating Source Information through Cross-Domain Dual-Branch Networks
III-C1 Training $M_{s}$ with Few Source Samples
@@
In the SF-UDA scenario with limited source domain samples, the primary challenge is how to learn semantic information from minimal data during source domain supervised training that is beneficial for unsupervised fine-tuning in the target domain. Traditional SF-UDA methods typically utilize a large number of samples to train model $M_{S}$, learning the feature distribution of the source domain. Subsequently, during the target domain training phase, information from the source domain is extracted from this model to construct a virtual domain. This surrogate aligns cross-domain feature distributions to enhance target domain performance, or further fine-tunes $M_{S}$ in the target domain through unsupervised or self-supervised methods to improve model performance.
However, in our scenario, the limited source domain samples make it challenging to acquire a complete feature distribution through supervised learning, let alone extract and align these features during target domain training. It is important to note that the domain disparity primarily stems from differences in image feature distributions, although the category names remain consistent between the source and target domains. Inspired by CLIP-based prompt learning techniques [56, 9], we propose extracting useful semantic information for the target domain by learning from category names in the sparse source domain samples. Upon completing the training, the text features of categories obtained through learning category names will be transmitted to the target domain, serving as a source of source domain category semantic information during target domain training.
@@
Specifically, as shown in the Figure 2, we freeze the parameters of CLIP and employ a commonly used textual prompt ”a photo of a” with $C$ learnable class-specific tokens $V^{L}=\{V_{0}^{L},V_{1}^{L}...,V_{C}^{L}\},V_{c}^{l}\in\mathcal{R}_{D}$, where $C$ denotes the number of classes. We concatenate the textual prompt with the class tokens and use CLIP’s text encoder $E_{text}$ to obtain the feature representation for each class as:
@@
where $E_{img}$ represents CLIP’s image encoder, and $\tau$ is a pre-defined temperature coefficient, set to 1 by default in this work. During the training of source domain, we optimize $V^{L}$ using the cross-entropy loss $\mathcal{L}_{\text{CE}}^{S}$ on the labeled source domain samples, as defined below:
@@
III-C2 Transferring Category Semantics Across Domains via Dual-Branch Networks
To fully leverage the category semantic information learned from source domain samples and to build a higher-performance model in the target domain, we propose a cross-modal dual-branch network. This network includes a cross-domain feature transfer branch aimed at obtaining predictions containing source domain category information, and a target-specific feature learning branch focused on acquiring target-specific predictions. Ultimately, we combine the outputs of both branches to form the model’s final prediction.
For the cross-domain feature transfer branch, inspired by prior work [54, 37], we propose enhancing the category text features learned from the source domain by incorporating a few image sample features that contain semantic information of target domain categories, thereby improving the classification ability of source domain text features for the target domain.
Specifically, as shown in Figure 2, we utilize the category feature representation $g^{S}$ learned from the source domain to compute the prediction results for all target domain samples using Equation 1. For each sample, the maximum category confidence is taken as its score, and the respective category is assigned as its pseudo-label. Next, based on the pseudo-labels, we partition $D_{T}$ into $C$ subsets $\{D_{T}^{1},D_{T}^{2},\dots,D_{T}^{C}\}$. For each subset, we select the $K$ samples with the highest category confidence as high-confidence samples. Finally, we aggregate the high-confidence samples from all subsets to form a highly reliable target domain sample set $D_{T}^{\prime}$. The reason for sorting within each subset is that the distribution of the maximum category confidence of CLIP’s zero-shot is inconsistent across categories [14], and sorting directly among all samples would result in a long-tail distribution of selected samples.

### v2

@@ Abstract
Source-free Unsupervised Domain Adaptation (SF-UDA) aims to transfer a model’s performance from a labeled source domain to an unlabeled target domain without direct access to source samples, addressing critical data privacy concerns. However, most existing SF-UDA approaches assume the availability of abundant source domain samples, which is often impractical due to the high cost of data annotation. To address the dual challenges of limited source data and privacy concerns, we introduce a data-efficient, CLIP-powered dual-branch network (CDBN). This architecture consists of a cross-domain feature transfer branch and a target-specific feature learning branch, leveraging high-confidence target domain samples to transfer text features of source domain categories while learning target-specific soft prompts. By fusing the outputs of both branches, our approach not only effectively transfers source domain category semantic information to the target domain but also reduces the negative impacts of noise and domain gaps during target training. Furthermore, we propose an unsupervised optimization strategy driven by accurate classification and diversity, preserving the classification capability learned from the source domain while generating more confident and diverse predictions in the target domain. CDBN achieves near state-of-the-art performance with far fewer source domain samples than existing methods across 31 transfer tasks on seven datasets.
@@
Source-Free, Unsupervised Domain Adaptation, Prompt Learning, CLIP, Data-Efficient
@@ I Introduction
Deep learning has achieved remarkable performance in various computer vision tasks [4, 59, 49]. However, these achievements largely rely on the assumption that the training and test data are independently and identically distributed (i.i.d.) [20]. In real-world scenarios, differences in data collection conditions, environmental factors, and other variables often cause domain shifts between the training data (source domain) and test data (target domain),significantly reducing model performance. To address this issue, previous works [25, 8] have proposed unsupervised domain adaptation (UDA), employing adversarial training [44, 8], metric learning [25], and other techniques to learn domain-invariant features, thereby facilitating the transfer of models trained on the source domain to the unlabeled target domain.
@@
Despite the effectiveness of UDA methods in addressing domain shifts, they typically assume that the source domain is accessible during target domain training and contains sufficient labeled data for training. However, this assumption is often unrealistic in practical settings due to the high cost of data annotation and the need to protect data privacy. For instance, in the diabetic retinopathy dataset used in clinical research [12], each image is annotated by 7 to 8 board-certified ophthalmologists, significantly increasing the cost of labeling [54]. Additionally, in fields such as medicine and remote sensing [36], labeled data often cannot be distributed or reused due to privacy protection and commercial copyright concerns.
@@
Figure 1: Scheme 1 illustrates the issue with previous SF-UDA methods, where using a fully trained model from a source domain as an information source for a target domain leads to poor performance when source domain samples are limited. To address this issue, Scheme 2 presents our proposed solution, which involves transferring category feature representations obtained from a source-domain-based prompt learning method to the target domain, effectively resolving the SF-UDA problem in scenarios with few source domain samples.
@@
Researchers have proposed two types of methods to address the challenges of restricted access to the source domain and the limited number of source domain samples. Source-free unsupervised domain adaptation (SF-UDA) methods [24, 44, 45] address the issue of inaccessible source data by enabling model transfer to the target domain without requiring direct access to the source domain. Meanwhile, few-shot unsupervised domain adaptation (FS-UDA) methods [54, 18, 23] tackle the problem of limited labeled data by facilitating effective model transfer with minimal source samples. However, methods that address the dual challenges of restricted access to source domain data and small amounts of source domain samples, known as SF-UDA methods under few-shot source domain scenarios (replaced by FS-SF-UDA in this paper), have yet to be fully explored.
@@
In recent years, prompt learning for CLIP models [17, 57, 10] has demonstrated significant advantages in few-shot classification tasks. These methods typically leverage CLIP as a backbone network and employ prompt learning techniques to achieve strong downstream task performance, even with only a limited number of samples per class. Inspired by this, we propose a data-efficient CLIP-powered Dual-Branch Network (dubbed as CDBN) for FS-SF-UDA.
@@
In scenarios with only a few samples in the source domain, our goal is to efficiently extract useful category semantic information from the source domain and deliver it to the target domain as a reference for unsupervised training. Specifically, as shown in Scheme 1 of Figure 1, traditional methods that rely on supervised training of a source domain model perform poorly when the number of samples in the source domain is limited. To address this issue, as illustrated in Scheme 2 of Figure 1, we employ a CLIP-based prompt learning approach to learn category text features with a certain degree of generalization from a small number of source domain samples. We then leverage these category text features as carriers of source domain semantic information to guide the unsupervised training process in the target domain.
@@
However, the combination of limited source domain samples and the domain gap presents a challenge, as it introduces noise and domain shift into the transferred category text features. To address this, we designed a cross-modal dual-branch network consisting of a cross-domain feature transfer branch and a target-specific feature learning branch. The cross-domain feature transfer branch utilizes high-confidence samples from the target domain to transfer the learned source domain category text features, reducing the impact of domain discrepancy. Meanwhile, the target-specific feature learning branch focuses on learning independent category features from the target domain, avoiding over-reliance on noisy source domain text features. The synergy between these two branches effectively enhances the feature representation capabilities in the target domain.
@@
Furthermore, since there are no labels during the training process in the target domain, we propose an unsupervised optimization strategy driven by accurate classification and diversity. This strategy includes three core loss functions: first, the cross-entropy loss from high-confidence target domain samples derived from source domain category text features ensures that the model retains the classification ability learned from the source domain; second, the multi-view consistency loss encourages the model to extract consistent feature representations from different perspectives, enhancing the generalization capability of the learned category feature distribution; finally, the mutual information maximization loss between inputs and outputs promotes diversity in the category distribution, preventing the model from overfitting to simple categories. Together, these three components ensure that the model can learn target domain features accurately and robustly under unsupervised conditions.
@@
We conducted experiments on 31 transfer tasks across 7 publicly available domain adaptation datasets. Despite the dual challenges of limited source samples and restricted access to source domain data, CDBN achieved state-of-the-art performance on multiple datasets and showed competitive results comparable to the best existing methods on others. These outcomes demonstrate the robustness and effectiveness of our CDBN approach in addressing the FS-SF-UDA problem.
@@
  • We introduce a cross-modal dual-branch network structure that effectively leverages source domain category semantic information during the unsupervised fine-tuning process in the target domain while mitigating the effects of source domain noise and domain discrepancy. Additionally, it combines the strengths of both branches, significantly enhancing performance in the target domain.
@@ II-B Few-Shot Unsupervised Domain Adaptation
Few-Shot Unsupervised Domain Adaptation (FS-UDA) aims to address the challenge of having limited labeled samples in the source domain during training on the target domain, even though source domain samples are accessible. Previous methods have explored different scenarios to tackle this issue. For instance, Yue et al. [54] proposes a method to address the problem of label sparsity in the source domain. In situations where a large number of source domain samples are available but only a small portion are labeled, they leveraged cross-domain self-supervised learning to adapt the model to the target domain. Similarly, Wang et al. [48] presents a method to handle the long-tailed distribution of source domain data. In cases where only a few samples are available for certain hard-to-access categories in the source domain, they improved model accuracy through generative feature augmentation. Furthermore, Li et al. [23] proposes a Multi-Prototype Alignment framework to transfer the model to the target domain when only a few samples per category are available in the source domain. This approach effectively addresses the scenario where there are very few samples for each category. In this paper, we tackle the challenges present in both FS-UDA and SF-UDA by training a high-performing model on the target domain despite having only a few samples per class in the source domain and no direct access to source domain samples during target domain training.
@@ II-C Prompt Learning for CLIP
In recent years, vision-language pretraining models have made significant progress in image representation learning by leveraging supervision from natural language to interpret images. Among these models, the Contrastive Language-Image Pre-training (CLIP [31]) model is particularly representative. CLIP consists of a text encoder and an image encoder, pre-trained on 400 million image-text pairs collected from the internet, successfully aligning the feature spaces of both modalities. This approach demonstrates strong generalization capabilities in downstream tasks and enables zero-shot transfer in image classification tasks through manually specified text prompts. Inspired by prompt learning techniques in natural language processing, Zhou et al. [57] introduces a learnable context optimization technique to adapt vision-language models like CLIP to image recognition tasks. Instead of relying on manually designed prompts, CoOp (Context Optimization) uses learnable vectors to model contextual words while keeping the pre-trained parameters frozen. This method allows CoOp to exhibit superior performance with limited labeled data, showing a significant improvement over manual prompt engineering in few-shot scenarios. To further enhance the alignment between visual and textual representations, Khattak et al. [17] proposes a method that applies prompt learning to both the image and text branches simultaneously, fostering a strong coupling between visual and language prompts to ensure their mutual synergy. As a result, Maple demonstrates significantly better generalization to novel classes, surpassing CoOp in performance.
Figure 2: The framework of CDBN, where $E_{img}$ represents the image encoder, $E_{text}$ represents the text encoder, $T_{t}$ denotes the class-shared soft prompt, and $V_{c}^{L}$ indicates the learnable token for the substitute class name of class c. Solid lines represent data flow, while dashed lines indicate the flow of values without gradient back-propagation. The flame symbol represents gradient updates, and the snowflake symbol indicates frozen gradients.
@@ III-A Problem Definition
Source-free unsupervised domain adaptation (SF-UDA) is an important sub-problem of unsupervised domain adaptation, focused on scenarios where source domain samples cannot be directly accessed during target domain training. This paper further explores an even more challenging scenario: the source domain samples are not only inaccessible but also limited in number.
Specifically, given a labeled source domain dataset $D_{S}=\{(x_{s}^{i},y_{s}^{i})\}_{i=1}^{n_{s}}$ with only a few samples and an unlabeled target domain dataset $D_{T}=\{x_{t}^{i}\}_{i=1}^{n_{t}}$, both $D_{S}$ and $D_{T}$ share the same labels but have different feature distributions. During the training of $D_{T}$, direct access to $D_{S}$ is not allowed. Instead, a model $M_{S}$ is first trained using $D_{S}$, and then $M_{S}$ is used to train on $D_{T}$. Finally, the model’s performance is evaluated on $D_{T}$.
@@ III-B Overview of CDBN
In this study, we aim to address the challenges of not directly accessing the source domain during target domain training, as well as the limited number of samples in the source domain. As shown in Figure 2, we propose a data-efficient CLIP-powered dual-branch network.
Specifically, we utilize the image encoder $E_{img}$ and text encoder $E_{text}$ from the CLIP model as the core network backbone. The model design includes a cross-domain feature transfer branch and a target-specific feature learning branch. The cross-domain feature transfer branch employs a cross-attention module with residual connections to fuse category text features $G^{s}$, obtained from the source domain through prompt learning, with the features $W_{e}$ of high-confidence pseudo-label samples in the target domain, facilitating feature transfer from the source domain to the target domain. The target-specific feature learning branch focuses on learning category-specific features for the target domain through a class-shared soft prompt mechanism,designed to minimize interference from noise in the source domain category text features. Finally, the model employs a fused logits network (FLN) and utilizes an unsupervised optimization strategy that combines the pseudo-label loss $L_{CE}^{T}$ from high-confidence target domain samples, multi-view consistency loss $L_{cons}$, and mutual information maximization loss $L_{IM}$ to train the model.
III-C Prompt Learning in the Source Domain
@@
In the SF-UDA scenario with limited source domain samples, the primary challenge is how to learn semantic information from minimal data during source domain supervised training that is beneficial for unsupervised fine-tuning in the target domain. Traditional SF-UDA methods typically utilize a large number of samples to train the model $M_{S}$, learning the feature distribution of the source domain. Subsequently, during the target domain training phase, information from the source domain is extracted from this model to construct a virtual domain. This surrogate aligns cross-domain feature distributions to enhance target domain performance or further fine-tunes $M_{S}$ in the target domain using unsupervised or self-supervised methods to improve model performance.
However, in our scenario, due to the limited number of samples in the source domain, it is challenging to obtain a complete feature distribution through supervised learning, let alone extract the source domain feature distribution for feature alignment during target domain training. In recent years, CLIP-based prompt learning methods [57, 10] have demonstrated the ability to learn feature representations with a certain degree of generalization from a small number of labeled samples. Inspired by these methods, we utilize a pre-trained CLIP model as the backbone of our model and learn category-specific, learnable tokens for the text encoder $E_{text}$. This enables us to obtain category text feature representations that have generalization capabilities for domain-shared class names from the limited source domain samples.
@@
Specifically, as shown in the Figure 2, we freeze the parameters of CLIP and employ a commonly used textual prompt ”a photo of a” with $C$ learnable class-specific tokens $V^{L}=\{V_{0}^{L},V_{1}^{L}...,V_{C}^{L}\},V_{c}^{l}\in\mathcal{R}_{D}$, where $C$ denotes the number of classes. We concatenate the textual prompt with the class-specific tokens and use CLIP’s text encoder $E_{text}$ to obtain the feature representation for each class as:
@@
where $E_{img}$ represents CLIP’s image encoder, and $\tau$ is a pre-defined temperature coefficient, set to 1 by default in this work. During the training of the source domain, we optimize $V^{L}$ using the cross-entropy loss $\mathcal{L}_{\text{CE}}^{S}$ on the labeled source domain samples, as defined below:
@@
III-D Cross-Modal Dual-Branch Networks
After obtaining the category text features learned from the source domain data, the limited sample size and domain discrepancy lead to noise and shifts in performance within the target domain. To retain useful category semantic information for the target domain while minimizing the introduction of noise that could affect the unsupervised training of target domain samples, we designed a cross-modal dual-branch network. This network includes a cross-domain feature transfer branch to adapt and correct for domain shifts, along with a target-specific feature learning branch to optimize the feature distribution specific to the target domain, preventing the model from overfitting to the category feature distribution learned from the source domain.
First, the function of the cross-domain feature transfer branch is to receive category text features from the source domain and transfer these features to the target domain using high-confidence target domain samples. This capability arises from the alignment of image and text feature distributions during CLIP’s pre-training, mapping them to a shared low-dimensional space, which allows image features to assist category text features in making predictions. Previous work [55] has demonstrated the feasibility of this approach by using a key-value store composed of labeled image features to enhance the classification of category text features obtained from manually designed text prompts. Inspired by earlier studies [1, 38], we propose to use a cross-attention module with residual connections to enable the source domain category text features to attend to the corresponding pseudo-labeled target domain image samples and facilitate their fusion, thereby transferring the source domain category text features to the target domain.
However, in our scenario, the target domain is unlabeled. To address this issue, we use the category text features from the source domain and the CLIP image encoder to obtain pseudo-labels for high-confidence samples in the target domain. As shown in the upper right corner of Figure 2, we utilize the category feature representation $g^{S}$ learned from the source domain to compute the prediction results for all target domain samples using Equation 1. For each sample, the maximum category confidence is taken as its score, and the respective category is assigned as its pseudo-label. Next, based on the pseudo-labels, we partition $D_{T}$ into $C$ subsets $\{D_{T}^{1},D_{T}^{2},\dots,D_{T}^{C}\}$. For each subset, we select the $K$ samples with the highest category confidence as high-confidence samples. Finally, we aggregate the high-confidence samples from all subsets to form a highly reliable target domain sample set $D_{T}^{\prime}$. The reason for sorting within each subset is that the distribution of the maximum category confidence of CLIP’s zero-shot is inconsistent across categories [15], and sorting directly among all samples would result in a long-tail distribution of selected samples.